# Data-driven Azimuthal RHEED construction for in-situ crystal growth characterization

### v3

@@ I Introduction
During MBE, it is essential to grow layers with substrate rotation to ensure uniformity Braun et al. (1998). Therefore, it is important to develop methods that allow access to the reciprocal lattice of rotating substrates Braun et al. (1998). Xiang et al. have showed in Xiang et al. (2016) that it is possible to obtain the entire reciprocal space structure of a 2D material by rotating the sample around the surface normal and measuring the RHEED patterns as a function of the azimuthal angle which makes ARHEED extremely interesting for surface characterization. A frequent challenge is the fact that a typical MBE chamber does not have a device for azimuthal angle measurement; and when it does, a typical difficulty is associated with the fact that in the ultra-high vacuum MBE chambers, the rotational stage is coupled magnetically and the rotation is often not perfectly smooth. This makes it difficult to determine the crystal angle with good accuracy. Determining the crystal azimuthal angle from the RHEED pattern itself is therefore very interesting for high-accuracy ARHEED. The Kikuchi lines, illustrated in Figure 4a-c with red lines, contain information on the sample orientation. Considering standard (001) oriented substrates, they are (for samples with a cubic crystal lattice) 4-fold symmetric and thus strictly contain information only for 0-90 degrees. Using several consecutive RHEED images instead of single ones to feed the neural network model would include information about the direction of the Kikuchi lines rotation and thus would allow to identify 0-180 degrees. Finally, the RHEED screen is usually not perfectly centered, such that each quadrant of the rotation shows a different section of the Kikuchi lines pattern, thus it should even be possible to distinguish between a full 0-360 degrees rotation angle range. Kikuchi lines originate from the electron scattering in the bulk of the crystal, hence do not (or only very weakly) depend on the growth of the topmost monolayer. Determination of the rotation angle should thus be possible in a robust way. Additionally, also the surface signal (diffraction spots and lines) contains information about the crystal angle, which can be used by a neural network as already demonstrated in earlier works Khaireh-Walieh et al. (2023); Kwoen and Arakawa (2020, 2022); Price et al. (2024). In summary, two elements are necessary in order to construct the ARHEED from raw RHEED patterns: the coordinates of the specular spot across the RHEED patterns and the rotational angle. Therefore, our approach in this article aims to construct the ARHEED from raw RHEED images by detecting the specular spot with segmentation using U-Net and determining the azimuthal angle via regression using ResNet. An overview of the proposed workflow is depicted in Figure 1.
@@ II.2.1 Dataset preprocessing
[…]
@@ II.3 Azimuthal RHEED plotting and interpretation
Figure 6: Surface reconstruction determination from azimuthal RHEED. Each azimuthal RHEED pattern contains information from two sample rotations. The large black segments correspond to the electron beam being blocked by the hooks of the sample holder. We set a threshold so that these parts appear dark black. This threshold sometimes weeds out other images that are initially almost black everywhere except at the specular point, this corresponds to the dark lines visible at certain angles. These patterns belong to a) $(2\times 4)$, b) transition and c) $c(4\times 4)$ surface reconstructions. The (00) point with the greatest intensity corresponds to the specular spot while the (01) and (10) spots correspond to the first order diffractions in [$\overline{1}$0] and [110] directions, respectively.
@@ II.4 Considerations about performance and potential real-time capabilities
Please note also, that the method as described here works only under the exact conditions that are captured by the training data. Our model does for example not work well on data acquired a few months after the training period, which we attribute to increased deposition on the view-port and variations in the electron beam configuration. We assume that the method can be extended to different growth environments using the fine-tuning technique. This enables a neural network model to adapt on new data that follows similar underlying correlations as the pre-training dataset. Fine-tuning requires less resources than training a model from scratch, yet the data should not be too different. We foresee that fine-tuning should work with similar RHEED systems (same acceleration voltage and similar distances), but full re-training may be necessary if the configurations are too different. It may also be possible to create a “foundation model” from fully synthetic data using computed Kikuchi line figures Pawlak et al. (2021b). Such model would not be useful for any actual application, but may be a versatile starting point for fine-tuning training on data from specific growth chambers.

### v4

@@ I Introduction
During MBE, it is essential to grow layers with substrate rotation to ensure uniformity Braun et al. (1998). Therefore, it is important to develop methods that allow access to the reciprocal lattice of rotating substrates Braun et al. (1998). Xiang et al. have showed in Xiang et al. (2016) that it is possible to obtain the entire reciprocal space structure of a 2D material by rotating the sample around the surface normal and measuring the RHEED patterns as a function of the azimuthal angle which makes ARHEED extremely interesting for surface characterization. A frequent challenge is the fact that a typical MBE chamber does not have a device for azimuthal angle measurement; and when it does, a typical difficulty is associated with the fact that in the ultra-high vacuum MBE chambers, the rotational stage is coupled magnetically and the rotation is often not perfectly smooth. This makes it difficult to determine the crystal angle with good accuracy. Determining the crystal azimuthal angle from the RHEED pattern itself is therefore very interesting for high-accuracy ARHEED. The Kikuchi lines, illustrated in Figure 4a-c with red lines, contain information on the sample orientation. Considering standard (001) oriented substrates, they are (for samples with a cubic crystal lattice) 4-fold symmetric and thus strictly contain information only for 0-90 degrees. Using several consecutive RHEED images instead of single ones to feed the neural network model would include information about the direction of the Kikuchi lines rotation and thus would allow to identify 0-180 degrees. Finally, the RHEED screen is usually not perfectly centered, such that each quadrant of the rotation shows a different section of the Kikuchi lines pattern, thus it should even be possible to distinguish between a full 0-360 degrees rotation angle range. Kikuchi lines originate from the electron scattering in the bulk of the crystal, hence do not (or only very weakly) depend on the growth of the topmost monolayer. Determination of the rotation angle should thus be possible in a robust way. Additionally, also the surface signal (diffraction spots and lines) contains information about the crystal angle, which can be used by a neural network as already demonstrated in earlier works Khaireh-Walieh et al. (2023); Kwoen and Arakawa (2020, 2022); pricePredictingAcceleratingNanomaterial2024. In summary, two elements are necessary in order to construct the ARHEED from raw RHEED patterns: the coordinates of the specular spot across the RHEED patterns and the rotational angle. Therefore, our approach in this article aims to construct the ARHEED from raw RHEED images by detecting the specular spot with segmentation using U-Net and determining the azimuthal angle via regression using ResNet. An overview of the proposed workflow is depicted in Figure 1.
@@ II.2.1 Dataset preprocessing
As explained before, the sample holder in the ultra high vacuum chamber is magnetically coupled to an external motor, often leading to non-smooth rotation. However, the sample holder has a significant physical weight. The faster we spin it, the less perturbation of the rotation by mechanical play and friction will occur thanks to the inertia of the sample holder. Therefore, angle accuracy is higher under fast spinning, but the ARHEED resolution reduces for high rotation speed. Good angle resolution requires rotation speeds in the order of 3-4 rpm, where friction and mechanical play have a very noticeable impact on the rotation. For this reason, we train a ResNet model for the azimuthal angle regression using a dataset of 26 300 images captured when the sample is rotating at 12 rpm for training. At this speed the rotation is very uniform and the relative angle can be accurately determined even without dedicated angular coder hardware, for example relative to a reference RHEED image. The dataset is split in such a way that 70 $\%$ is used for the model training, 15 $\%$ for validation and 15 $\%$ for test. For the ARHEED construction, the test dataset of the specular beam tracking in the above section is used. The raw test images have $(512\times 688)$ dimension in 16-bit (65536 grayscale) and are captured at 4 rpm. In preprocessing, we reduce the image dimensions to $(137\times 229)$ and at the same time normalize pixel values between 0 and 1. After tests of different size scale factors, we choose the smallest dimensions without noticeable impact on the results. Sequences of 6 images bring better accuracy for the regression task thus the model input dimensions are $(137\times 229\times 6)$, stacking six consecutive RHEED images along the channel dimension. We define the prediction angle to be the angle corresponding to the first image of each sequence. For a real-time application, the oldest, last image would be removed from the stack and the newest image, obtained from the RHEED camera, would be added as new first image for each prediction. We use the open source software Kikuchipy Ånes et al. (2020) for RHEED image pre-processing, which allows to increase the contrast of the Kikuchi lines, thereby enabling the model to focus on their movement. We found that this pre-processing significantly improves the training convergence of the azimuthal angle prediction network. Kikuchi lines depend mainly on the bulk of the crystal Pawlak et al. (2021), which leads us to believe that this is a robust method for deducing the rotation angle from RHEED images. Two pre-processing steps are applied on the images using Kikuchipy: removing of dynamic background and adaptive histogram equalization, as illustrated in Figure 4b and 4c, respectively. Figure 4c shows that the contrast of the Kikuchi lines is signficiantly increased after this pre-processing. Note, that the neural network is trained on the full RHEED images, apart from the Kikuchi lines contrast pre-processing, we implemented no further techniques that would force the models to use exclusively Kikuchi lines for angular regression. To study the impact of Kikuchi visibility or signal-to-noise ration, synthetic data would be required, which would offer the possibility to arbitrarily tune the Kikuchi line contrast or the amount of noise only on the Kikuchi lines.
@@ II.3 Azimuthal RHEED plotting and interpretation
[…]
@@ II.4 Considerations about performance and potential real-time capabilities
Please note also, that the method as described here works only under the exact conditions that are captured by the training data. Our model does for example not work well on data acquired a few months after the training period, which we attribute to increased deposition on the view-port and variations in the electron beam configuration. We assume that the method can be extended to different growth environments using the fine-tuning technique. This enables a neural network model to adapt on new data that follows similar underlying correlations as the pre-training dataset. Fine-tuning requires less resources than training a model from scratch, yet the data should not be too different. We foresee that fine-tuning should work with similar RHEED systems (same acceleration voltage and similar distances), but full re-training may be necessary if the configurations are too different. It may also be possible to create a “foundation model” from fully synthetic data using computed Kikuchi line figures pawlakAnalysisKikuchiLines2021. Such model would not be useful for any actual application, but may be a versatile starting point for fine-tuning training on data from specific growth chambers.